Which of the following is not an equal set?
i) A = {4, 3, 2, 1 } B= {1, 3, 2, 4 }
ii) A ={ 2, 3 } , B = { 3, 2 }, C = { x:x2 – 5x – 6 =0 }
iii) A = { 4, 5 ) B = { 5, 4 }
iv) A = { 6, 7, 8 } B= { Six, seven, eight }

Answer by Edwin McCravy(20055)

Which of the following is not an equal set?
i) A = {4, 3, 2, 1 } B= {1, 3, 2, 4 }
Those are equal sets because the order in which they are written
does not matter.

ii) A ={ 2, 3 } , B = { 3, 2 }, C = { x:x² – 5x – 6 =0 }
A and B are equal sets because the order in which elements are written 
does not matter. Let's solve the equation in C to find out what
elements it has:

  x² - 5x - 6 = 0
   (x+1)(x-6) = 0
  x+1 = 0;  x-6 = 0
    x = -1;   x = 6

Aha!  That's the set C = {-1, 6}, not the set (2, 3}.  

[Note:  If it had been C = { x:x² – 5x + 6 =0 }, with +6 instead 
of -6, C would have been the same set as A and B]

iii) A = { 4, 5 ) B = { 5, 4 }
Those are equal sets because the order in which they are written does not matter.

iv) A = { 6, 7, 8 } B= { Six, seven, eight }
Those are equal sets because the numerals for the numbers and the words
for the numbers are both symbols for the same numbers.

Edwin